The sum of the digits is 12.
The digit in the tousandths place is 3 times the digit in the tens place.
To the nearest ten, the number rounds to 30.
The digit in the tenths place is one more than the digit in the hundredths place.
What is the number?

Answer by ankor@dixie-net.com(22740)

The three digits: a, b, c, d
:
"The sum of the digits is 12."
a + b + c + d = 12
:
"The digit in the thousandths place is 3 times the digit in the tens place."
a = 3c
"The digit in the tenths place is one more than the digit in the hundredths place."
c = b + 1
or
b = (c-1)
:
a + b + c + d = 12
Replace a and b
3c + (c-1) + c + d = 12
5c + d = 12 + 1
5c + d = 13
Assume that c has to equal 2, therefore
5(2) + d = 13
10 + d = 13
d = 13 - 10
d = 3
find the rest of the digits
a = 3(2) = 6
b = 2-1 = 1
c = 2
d = 3
then
6123 is the number
:
Assume that c has to equal 1, therefore
5(1) + d = 13
5 + d = 13
d = 13 - 5
d = 8
find the rest of the digits
a = 3(1) = 3
b = 1-1 = 0
c = 1
d = 8
then
3018 is the number
:
"To the nearest ten, the number rounds to 30." Not sure what bearing this has on this problem.